OpenStudy (anonymous):

A bit confused. So, I am trying to put (1-i)^3 into standard form, but when I do (1-i)*(1-i) I get zero.

OpenStudy (sgadi):

I dont udnerstand why (1-i)(1-i)1=0. i propose (1-i)(1-i)(1-i)=(1-1+2)(1-i)=2-2i